# skipabove=1em,skipbelow=1em

### v3

@@ 4. Taxonomy of Target Factors and Behaviors
Daily ingestive behavior interventions can play an important role in promoting healthier eating and drinking habits. The targets of these interventions reflect the key behaviors believed to be critical for promoting healthy eating and improving overall health outcomes. Understanding the specific factors and behaviors targeted in the existing intervention literature is crucial, as they not only highlight important aspects of ingestive behaviors relevant to health, but also provide insight into the intervention strategies designed to address them. By categorizing and analyzing these targeted behaviors, we can better understand how current interventions aim to modify eating habits, assess the effectiveness of different approaches, and identify gaps in existing methods.
This chapter aims to explore the targeted behaviors in current ingestive behavior intervention studies, categorized into contextual factors (external to the human body) and process behaviors (individual actions during the ingestion process). These two broad categories encompass a wide range of interventions designed to optimize various target factors or behaviors for ingestion and promote healthier ingestion habits.
Figure 3 presents trends in the number of intervention studies over the years which can be categorized into three groups: contextual factors only, process behaviors only, and studies targeting both. Both categories have garnered significant attention from researchers, with a greater number of interventions focusing on contextual factors.
@@
Building on this, Figure 6 displays the taxonomy of these target factors and behaviors, providing a structured overview of the specific intervention targets within each category and the number of studies associated with each. This taxonomy organizes the diverse intervention targets into a hierarchical structure, clarifying the relationship between broader categories and more specific target behaviors. A summary of these categories and the corresponding publications included in this review is shown in Table 6.
@@ 4.1. Contextual Factors
Contextual factors refer to external influences that affect ingestive behaviors, which are usually not limited to the duration of the ingestion process itself. These factors, which include food choices, intake timing, the environment in which eating takes place, and activity around ingestion process, can shape an individual’s dietary patterns and habits. Since these influences often operate outside the immediate context of the meal, they play a crucial role in determining overall ingestive behaviors and are frequently targeted in interventions aimed at promoting healthier eating habits.
@@ 4.1.1. Food Choices
Given the central role of food choices in overall dietary health, many current ingestive behavior interventions prioritize food choices as a primary target. In fact, among the reviewed literature, interventions targeting food choices constitute the largest proportion. These interventions aim to improve dietary quality by modifying nutrient intake, promoting healthier eating habits, and restricting the consumption of unhealthy foods.
Within these interventions, food choice-related targets are commonly categorized into several specific areas: Nutrition Structure Control (aiming for balanced nutrition and dietary variety), Calorie Control (regulating total energy intake), Macronutrient and Micronutrient Control (focusing on specific nutrients such as carbohydrates, fats, sugars, and salt), Limiting Specific Food Categories (e.g., reducing intake of snacks and sugary beverages), and Unspecified Food Choices (general recommendations for healthier eating without specific nutrient guidelines). Among these categories, interventions focusing on nutrition balance and calorie control are particularly prevalent, reflecting their perceived critical role in promoting healthy eating and preventing diet-related chronic diseases.
Nutrition Structure Control: This target behavior focuses on ensuring balanced nutrient intake and promoting dietary variety.
Nutrition structure control interventions focus on fostering a balanced intake of various nutrients by encouraging dietary diversity. Rather than emphasizing individual food items or nutrients, these interventions encourage individuals to consume a wide range of food groups, including fruits, vegetables, whole grains, proteins, and dairy. A significant portion of these interventions specifically target children, with the goal of instilling balanced eating habits early in life. Many studies aim to prevent picky ingestive behaviors and encourage the development of healthy, varied eating patterns from a young age. By promoting a broader range of food intake, these interventions help address nutrient deficiencies and support long-term health maintenance.
Calorie Target Control: This target behavior aims to regulate energy intake through portion size estimation or meal planning.
Calorie control interventions primarily aim to manage overall energy intake to prevent overweight, obesity, and related chronic diseases such as diabetes and cardiovascular conditions. These interventions guide individuals in selecting lower-calorie, nutrient-dense foods, such as fruits, vegetables, and lean proteins. Strategies often include reducing portion sizes and replacing high-calorie foods with healthier alternatives. In the reviewed literature, calorie control is frequently employed as a strategy for weight management and reducing obesity risk, underscoring its central role in ingestion health interventions.
Macronutrient and Micronutrient Control: This category includes interventions targeting the regulation of specific nutrient intakes, such as carbohydrates, fats, sugars, and salt. Many interventions focus on adjusting these components to address health issues like diabetes, hypertension, and cardiovascular diseases. For example, regulating carbohydrate intake is often emphasized for individuals at risk of or diagnosed with diabetes, while reducing fat (particularly saturated fat) and sodium intake is associated with cardiovascular risk reduction.
@@
  • Salt Intake — Limiting sodium intake helps prevent hypertension and cardiovascular risks (Jaques et al., 2021). Current interventions focus on controlling salt intake to promote healthier eating habits in patients with metabolic syndrome, such as those with cardiovascular disease (Eyles et al., 2017; Kim et al., 2016c).
Limiting Specific Food Categories: Some interventions specifically target the reduction or elimination of certain food categories considered unhealthy, such as snacks high in sugar or fat, and sugary beverages like soft drinks and alcohol. By limiting the intake of these foods, the aim is to reduce empty calorie consumption, lower the risk of obesity, and improve overall diet quality. This approach often complements broader nutrition and calorie-focused strategies by removing key sources of dietary excess.
@@
Unspecific Food Choices: This refers to general recommendations that do not focus on specific nutrients or food categories.
A smaller subset of interventions addresses food choices in a more general way, providing broad guidance to ”eat healthier” without specifying particular nutrients or food categories. These interventions typically focus on encouraging awareness about healthy eating principles and promoting better food decision-making habits, sometimes through educational tools (Prasetyo et al., 2020; Fernandez et al., 2013), reminders (Abisha and Rajalakshmy, 2017; Luo et al., 2021), or behavioral nudges (Huang et al., 2018; Terzioğlu et al., 2023; Eigen et al., 2018).
@@ 4.1.2. Intake Timing
Intake timing refers to the timing factors of ingestion, which significantly influence metabolic processes, health, and satiety. Studies have consistently shown that meal timing and frequency are closely linked to overall metabolic health, including insulin sensitivity (Alkhulaifi and Darkoh, 2022). Unhealthy eating patterns, such as irregular meal timing or late-night eating, have been associated with a range of adverse health outcomes. Disruptions to the natural timing of meals can contribute to the development of chronic diseases, including obesity, cardiovascular disease, and diabetes (Yoshida et al., 2018; Rask-Madsen and Kahn, 2012). Additionally, meal frequency — how often a person eats throughout the day — has been found to influence metabolic health. Appropriate meal frequency may promote weight management and improve insulin sensitivity and blood glucose levels (Holmbäck et al., 2010; Carlson et al., 2007). Along with maintaining regular meal times, consistent fluid intake is also essential for supporting bodily functions, such as temperature regulation, nutrient transportation, and waste removal.
@@ 4.1.3. Environment
The environment in which an individual eats can significantly impact ingestive behaviors. Both social, cultural, and physical environments contribute to shaping food choices, eating patterns, and the overall eating experience. For example, the availability of healthy food options, social influences, and food cues in the surroundings can either promote or hinder healthy ingestive behaviors. Research has shown that eating location and the surrounding environment can notably affect the intake of unhealthy foods. For instance, higher intakes of unhealthy snacks, such as those rich in sugars, fats, and sodium, are often observed in less structured eating environments like schools and while socializing (Casey et al., 2021).
@@ 4.2. Process Behaviors
Process behaviors refer to the specific actions individuals perform during ingestion. Interventions targeting process behaviors typically occur within the duration of an ingestion episode, although they may also span a longer ingestion period. These behaviors, including intake quantity, ingestion pace and oral motor activities, food awareness or emotional state, and safe eating practices, directly impact food intake and digestion. argeting these aspects aims to improve eating habits, prevent overeating, promote mindful consumption, and reduce eating-related safety risks.
@@ 4.2.1. Intake Quantity
Solid Food: In this context, solid food refers to the ingestion of meals or snacks, excluding water and other beverages. By promoting better awareness and control of food quantities, these interventions aim to prevent overeating, reduce the risk of obesity, and support healthy ingestive behaviors, including the prevention of eating disorders (ED) (Rolls et al., 2002; Livingstone and Pourshahidi, 2014).
Interventions targeting solid food portion control often measure intake quantity by assessing changes in food mass before and after a meal (Xie et al., 2023; Chaudhry et al., 2016; Chen et al., 2022; Alshurafa et al., 2015; Scisco et al., 2011; Maramis et al., 2014; Farooq et al., 2017; Moses et al., 2023; Li and Bailenson, 2018).
Liquid: In this category, liquid intake specifically refers to the consumption of water and beverages. Most interventions targeting liquid intake emphasize the importance of maintaining adequate hydration throughout the day. Sufficient fluid intake is crucial for supporting various physiological functions, including temperature regulation, cardiovascular and cognitive health, and preventing conditions such as kidney stones and urinary tract infections. It may also aid in managing energy intake and weight by promoting healthier beverage choices (Popkin et al., 2010; Daniels and Popkin, 2010).
Interventions focusing on fluid intake typically measure water consumption by tracking changes in volume, such as the amount consumed between two time points or the total daily intake (Faltaous et al., 2021; Alshurafa et al., 2015; Kaner et al., 2018; Poddar et al., 2024; Kreutzer et al., 2015; Varun et al., 2023; Bobin et al., 2018; Lessel et al., 2016; Yıldız and Coşkun, 2019; Chiu et al., 2009; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019; Tommy et al., 2017; Ravindran et al., 2022; Zhou et al., 2021).
@@ 4.2.2. Ingestion Pace and Oral Motor Activities
Ingestion pace and oral motor behaviors play a crucial role in regulating energy intake and promoting healthy eating patterns. Interventions targeting these behaviors focus on modifying the speed and manner in which food is consumed to enhance satiety, prevent overeating, and support weight management. Faster eating rates have been consistently associated with higher energy intake, increased risk of overweight and obesity, and poorer appetite regulation. A growing body of evidence supports the effectiveness of slowing down ingestion pace as a practical strategy for preventing obesity and related chronic conditions (Robinson et al., 2014; Slyper, 2021). Current approaches typically target three main aspects: chewing and swallowing behaviors, bite intervals and size, and total meal duration.
@@
Chewing and swallowing behaviors: Chewing and swallowing behaviors are key components influencing the pace of eating and the overall ingestion process. Many interventions target chewing behavior specifically by encouraging individuals to chew more times per bite or to slow down the rate of chewing. By increasing the number of chews per mouthful, individuals can naturally slow their eating pace, enhance satiety signaling, and reduce overall food intake, which are important strategies for managing weight and preventing overeating (Kadomura et al., 2013b; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Chen and Yen, 2022; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023).
Some interventions combine the regulation of chewing and swallowing, aiming to control the full oral processing cycle rather than focusing solely on chewing. These approaches aim to slow ingestion by coordinating the timing between chewing and swallowing, ensuring thorough oral processing before taking the next bite (Kim et al., 2016b; Kamachi et al., 2023).
Additionally, a few interventions focus on modifying oral motor behaviors by adjusting chewing force and modulating the swallowing phase. These interventions promote safer, more mindful eating practices, improving digestion and reducing the risk of eating-related complications (Biyani et al., 2019; Zheng et al., 2025).
@@
Bite Intervals and Size: Bite size and the interval between bites are critical aspects of eating pace regulation and have been targeted in current interventions. This strategy allows more time for satiety signals to develop, helping individuals recognize fullness earlier and thus reducing overall food intake. Most interventions focus on extending the time interval between successive bites, encouraging individuals to pause longer before taking the next bite (Nakamura et al., 2023; de Vries et al., 2020; Lee et al., 2019; Kim et al., 2016a; Kim and Bae, 2018b, a; Mendi et al., 2013; Hermans et al., 2017). In addition to manipulating bite intervals, the amount of food consumed in a single bite is also a key target behavior. By reducing bite size, individuals naturally slow down their eating pace and consume fewer calories during a meal (Chen et al., 2022).
@@
Total Meal Duration: Total meal duration — the overall time taken to complete a meal — is another important target for interventions aimed at regulating eating pace. Most studies monitor and estimate total meal duration by tracking changes in food weight over time, providing an objective measure of ingestion pace. Many interventions focus on prolonging meal duration to slow down eating speed, helping individuals better recognize satiety cues and prevent overeating, which supports healthier dietary behaviors (Moses et al., 2023; Zhang et al., 2019; Kim et al., 2016c; Ioakimidis et al., 2009). Some interventions, particularly those targeting young children, aim to prevent meals from extending excessively long. These interventions help children establish structured, healthy eating routines early in life (Krishna et al., 2025; Lo et al., 2007). Additionally, some interventions are designed to synchronize eating paces in social settings, where regulating total meal duration across a group helps reduce discomfort caused by uneven eating speeds (Mitchell et al., 2015).
@@ 4.2.3. Food Awareness or Emotional State
Interventions targeting food awareness or emotional state focus on enhancing individuals’ cognitive and emotional engagement with their eating experiences to promote healthier behaviors. A key approach in this area is encouraging greater attention to food during meals, either through practices like mindful eating, which emphasizes being fully present during eating (Kleinberger et al., 2023; Koizumi et al., 2011; Parra et al., 2023; Wang et al., 2025; Khot et al., 2020), or through interventions designed to help children develop better focus on their food (Kadomura et al., 2013b; Han and Kang, 2017). In addition to promoting awareness during ingestion, some interventions focus on improving individuals’ self-perception and cognitive evaluation of their ingestive behaviors after meals. By increasing reflective thinking about eating habits, these interventions help individuals build healthier long-term dietary patterns and strengthen their self-regulation capacity (Sun et al., 2020; Zhang and Parker, 2020; Tang et al., 2015). Moreover, another line of interventions aims to enhance the emotional quality of the eating experience itself, such as promoting positive emotions through novel experiences and reducing negative emotions like anxiety during eating. A healthy emotional state during ingestion has been linked to better digestion, improved dietary choices, and a lower risk of disordered ingestive behaviors (Mayumi et al., 2023; Carroll et al., 2013; Mitchell et al., 2015; Narumi et al., 2011).

### v4

@@ 4. Taxonomy of Target Factors and Behaviors
Daily ingestive behavior interventions can play an important role in promoting healthier eating and drinking habits. Identifying the specific factors and behaviors targeted in prior research helps clarify health-relevant aspects of ingestion and associated intervention strategies. By analyzing these behaviors, we can evaluate how current interventions modify eating patterns, compare approach effectiveness, and uncover gaps in existing methods.
This chapter aims to explore the targeted behaviors in current ingestive behavior intervention studies, and classify them into contextual factors (external to the human body) and process behaviors (individual actions during the ingestion process). These two broad categories encompass a wide range of interventions designed to optimize various target factors or behaviors for ingestion and promote healthier ingestion habits.
Figure 3 presents the yearly distribution of intervention studies which can be categorized into three groups: contextual factors only, process behaviors only, and studies targeting both. Both categories have garnered significant attention from researchers, with a greater number of interventions focusing on contextual factors. Building on this, Figure 6 displays the taxonomy of these target factors and behaviors, providing a structured overview of the specific intervention targets within each category and the number of associated studies. This taxonomy organizes the diverse intervention targets into a hierarchical structure, clarifying the relationship between broader categories and more specific target behaviors. A summary of these categories and the corresponding publications included in this review is shown in Table 2.
@@ 4.1. Contextual Factors
Contextual factors refer to external influences that affect ingestive behaviors, which are usually not limited to the duration of the ingestion process itself. These include food choices, intake timing, the eating environment, and activities surrounding the ingestion process, all of which can shape an individual’s dietary patterns and habits.
@@ 4.1.1. Food Choices
In fact, interventions focusing on food choices represent the largest portion of the reviewed literature. These interventions aim to improve dietary quality by modifying nutrient intake, promoting healthier eating habits, and restricting the consumption of unhealthy foods. Within these interventions, food choice-related targets are commonly categorized into several specific areas: Nutrition Structure Control, Calorie Control, Macronutrient and Micronutrient Control, Limiting Specific Food Categories, and Unspecified Food Choices. Among these categories, interventions focusing on nutrition balance and calorie control are particularly prevalent, reflecting their perceived critical role in promoting healthy ingestion and preventing diet-related chronic diseases.
i. Nutrition Structure Control: This target behavior aims to promote balanced nutrient intake by encouraging dietary diversity across major food groups, such as fruits, vegetables, whole grains, proteins, and dairy. A significant portion of interventions targets children, seeking to cultivate healthy eating habits early in life. Additionally, some studies address picky eating behaviors that limit dietary variety. By broadening food intake, these interventions help prevent nutrient deficiencies and support long-term health maintenance.
ii. Calorie Target Control: This target behavior focuses on regulating energy intake through portion control, meal planning, and calorie management. Interventions guide individuals toward lower-calorie, nutrient-dense food, such as fruits, vegetables, and lean proteins, while encouraging the replacement of high-calorie items with healthier alternatives. In the reviewed literature, calorie control is frequently employed as a strategy for weight management and obesity prevention, underscoring its central role in ingestion health interventions.
iii. Macronutrient and Micronutrient Control: This category includes interventions targeting the regulation of specific nutrient intakes, such as carbohydrates, fats, sugars, and salt. Many interventions focus on adjusting these components to address health issues like diabetes, hypertension, and cardiovascular diseases.
@@
  • Salt Intake — Limiting sodium intake helps prevent hypertension and cardiovascular risks (Jaques et al., 2021). Current interventions focus on controlling salt intake in patients with metabolic syndrome, such as those with cardiovascular disease (Eyles et al., 2017; Kim et al., 2016c).
iv. Limiting Specific Food Categories: Some interventions specifically target the reduction or elimination of certain food categories considered unhealthy, such as snacks high in sugar or fat, and sugary beverages like soft drinks and alcohol.
@@
v. Unspecific Food Choices: This refers to general recommendations that do not focus on specific nutrients or food categories. A smaller subset of interventions addresses food choices in a more general way, providing guidance to ”eat healthier” without specifying particular nutrients or food categories. These interventions typically focus on encouraging awareness about healthy eating principles and promoting better food decision-making habits, sometimes through educational tools (Prasetyo et al., 2020; Fernandez et al., 2013), reminders (Abisha and Rajalakshmy, 2017; Luo et al., 2021), or behavioral nudges (Huang et al., 2018; Terzioğlu et al., 2023; Eigen et al., 2018).
@@ 4.1.2. Intake Timing
Intake timing refers to the timing factors of ingestion, which significantly influence metabolic processes, health, and satiety. Studies have consistently shown that meal timing and frequency are closely linked to overall metabolic health, including insulin sensitivity (Alkhulaifi and Darkoh, 2022). Unhealthy eating patterns, such as irregular meal timing or late-night eating, have been associated with a range of adverse health outcomes. Disruptions to the natural timing of meals can contribute to the development of chronic diseases, including obesity, cardiovascular disease, and diabetes (Yoshida et al., 2018; Rask-Madsen and Kahn, 2012). Additionally, meal frequency — how often a person eats throughout the day — has been found to influence metabolic health. Appropriate meal frequency may promote weight management and improve insulin sensitivity and blood glucose levels (Holmbäck et al., 2010; Carlson et al., 2007).
@@ 4.1.3. Environment
Ingestive behaviors are shaped by a combination of social, cultural, and physical environments, which influence food choices, eating patterns, and the broader eating experience. Research has shown that eating location and the surrounding environment can notably affect the intake of unhealthy foods. For instance, higher intakes of unhealthy snacks, such as those rich in sugars, fats, and sodium, are often observed in less structured eating environments like schools and while socializing (Casey et al., 2021).
@@ 4.2. Process Behaviors
Process behaviors refer to the specific actions individuals perform during ingestion. Interventions targeting process behaviors typically occur within the duration of an ingestion episode, although they may also span a longer ingestion period. These behaviors-including intake quantity, ingestion pace, oral motor patterns, food awareness, emotional state, and safe eating practices—directly influence food intake and digestion.
@@ 4.2.1. Intake Quantity
i. Solid Food: Solid food refers to the ingestion of meals or snacks, excluding water and other beverages. These interventions aim to enhance awareness and control of food quantities to prevent overeating and reduce obesity risk. They also support healthy ingestive behaviors, including the prevention of eating disorders (Rolls et al., 2002; Livingstone and Pourshahidi, 2014). Interventions targeting solid food portion control often measure intake quantity by assessing changes in food mass before and after a meal (Xie et al., 2023; Chaudhry et al., 2016; Chen et al., 2022; Alshurafa et al., 2015; Scisco et al., 2011; Maramis et al., 2014; Farooq et al., 2017; Moses et al., 2023; Li and Bailenson, 2018).
ii. Liquid: Liquid intake refers to the consumption of water and beverages. Most interventions in this category emphasize maintaining adequate hydration to support physiological functions and promote healthier beverage choices. These strategies may also aid in energy balance and weight management (Popkin et al., 2010; Daniels and Popkin, 2010). Interventions focusing on fluid intake typically measure water consumption by tracking changes in volume, such as the amount consumed between two time points or the total daily intake (Faltaous et al., 2021; Alshurafa et al., 2015; Kaner et al., 2018; Poddar et al., 2024; Kreutzer et al., 2015; Varun et al., 2023; Bobin et al., 2018; Lessel et al., 2016; Yıldız and Coşkun, 2019; Chiu et al., 2009; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019; Tommy et al., 2017; Ravindran et al., 2022; Zhou et al., 2021).
@@ 4.2.2. Ingestion Pace and Oral Motor Activities
Interventions targeting ingestion pace and oral motor behaviors focus on modifying the speed and manner in which food is consumed to enhance satiety, prevent overeating, and support weight management. Faster eating rates have been consistently associated with higher energy intake, increased risk of overweight and obesity, and poorer appetite regulation. A growing body of evidence supports the effectiveness of slowing down ingestion pace as a practical strategy for preventing obesity and related chronic conditions (Robinson et al., 2014; Slyper, 2021). Current approaches typically target three main aspects: chewing and swallowing behaviors, bite intervals and size, and total meal duration.
@@
i. Chewing and swallowing behaviors: Chewing and swallowing behaviors are key components influencing the pace of eating and the overall ingestion process. Many interventions encourage individuals to chew more times per bite or slow down the rate of chewing. These practices can naturally slow their eating pace, enhance satiety signaling, and reduce overall food intake, which supports weight management and prevention of overeating (Kadomura et al., 2013b; Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Scisco et al., 2011; Maramis et al., 2014; Chen and Yen, 2022; Sugita et al., 2018; Turner-McGrievy et al., 2017; Hori et al., 2023). Others combine the regulation of chewing and swallowing, aiming to control the full oral processing cycle rather than focusing solely on chewing. These approaches aim to slow ingestion by coordinating the timing between chewing and swallowing, ensuring thorough oral processing before taking the next bite (Kim et al., 2016b; Kamachi et al., 2023). Additionally, a few interventions focus on modifying oral motor behaviors by adjusting chewing force and modulating the swallowing phase. These interventions promote safer, more mindful eating practices, improving digestion and reducing the risk of eating-related complications (Biyani et al., 2019; Zheng et al., 2025).
@@
ii. Bite Intervals and Size: As critical components of eating pace regulation, bite size and inter-bite intervals have been intervened to allow more time for satiety signals to develop, helping individuals recognize fullness earlier and thus reducing overall food intake. Most interventions focus on extending the time interval between successive bites, encouraging individuals to pause longer before taking the next bite (Nakamura et al., 2023; de Vries et al., 2020; Lee et al., 2019; Kim et al., 2016a; Kim and Bae, 2018b, a; Mendi et al., 2013; Hermans et al., 2017). Bite size is another key behavioral target. Reducing the amount of food per bite naturally slows eating pace and lowers overall caloric intake during a meal (Chen et al., 2022).
@@
iii. Total Meal Duration: The overall time taken to complete a meal is typically estimated by tracking changes in food weight over time to provide an objective measure of ingestion pace. Many interventions focus on prolonging meal duration to slow down eating speed, helping individuals better recognize satiety cues and prevent overeating (Moses et al., 2023; Zhang et al., 2019; Kim et al., 2016c; Ioakimidis et al., 2009). Others, particularly those involving young children, seek to avoid excessively long meals to help establish structured eating routines early in life (Krishna et al., 2025; Lo et al., 2007). Additionally, some interventions are designed to synchronize eating paces in social settings, where regulating total meal duration across a group helps reduce discomfort caused by uneven eating speeds (Mitchell et al., 2015).
@@ 4.2.3. Food Awareness or Emotional State
Interventions targeting food awareness or emotional state focus on enhancing individuals’ cognitive and emotional engagement with their eating experiences to promote healthier behaviors. A key approach is encouraging greater attention to food during meals, either through practices like mindful eating, which emphasizes being fully present during eating (Kleinberger et al., 2023; Koizumi et al., 2011; Parra et al., 2023; Wang et al., 2025; Khot et al., 2020), or through interventions designed to help children develop better focus on their food (Kadomura et al., 2013b; Han and Kang, 2017). Other interventions focus on improving individuals’ self-perception and cognitive evaluation of their ingestive behaviors after meals. By increasing reflective thinking about eating habits, these interventions help individuals build healthier long-term dietary patterns and strengthen their self-regulation capacity (Sun et al., 2020; Zhang and Parker, 2020; Tang et al., 2015). Moreover, a separate line of interventions aims to enhance the emotional quality of the eating experience itself, such as promoting positive emotions through novel experiences and reducing negative emotions like anxiety during eating. A healthy emotional state during ingestion has been linked to better digestion, improved dietary choices, and a lower risk of disordered ingestive behaviors (Mayumi et al., 2023; Carroll et al., 2013; Mitchell et al., 2015; Narumi et al., 2011).